OpenStudy (anonymous):

PLEASE HELP !!! Jeff is baking a cake. The recipe says that he has to mix 32 grams of vanilla powder to the flour. Jeff knows that 1 cup of that particular vanilla powder has a mass of 128 grams. He added two over three of a cup of vanilla powder to the flour. Should Jeff add more vanilla powder to make the exact recipe or did he go over and by what amount?

OpenStudy (anonymous):

summary of the slippery slope by lemony snicket if anyone knows

OpenStudy (johnweldon1993):

Well he needed 32 grams What he added...was 2/3 a cup ... we know 1 cup = 128 grams...so to figure out how much he added...we multiply 128 by 2/3 \[\large 128 \times \frac{2}{3}\] \[\large \frac{128 \times 2}{3}\] \[\large \frac{256}{3} = \space ?\]